OpenStudy (anonymous):

if sin theta -cos theta=0 then sin(pi/2-theta)+cos(pi/2-theta)=?

OpenStudy (anonymous):

cos theta+sin theta=

OpenStudy (ikram002p):

sin theta -cos theta=0 then tan theta =1 theta = tan^-1 (1) so theta = ?

OpenStudy (ikram002p):

or you wanna you sum formula ?

OpenStudy (anonymous):

i need the value of sin theta+cos theta=?

OpenStudy (ikram002p):

well theta = pi/4 :o

OpenStudy (ikram002p):

@ganeshie8 any other way to solve it ?

OpenStudy (anonymous):

can we do like that sin pi/2-sin theta +cos pi/2-cos theta?

OpenStudy (ikram002p):

well cos theta =sin theta and we wanna find cos theta +sin theta =2 cos theta =2 sin theta xD which i have no idea what is this xD

ganeshie8 (ganeshie8):

we can simply substitute \(\theta = \pi/4\) right ?

OpenStudy (anonymous):

2 sin pi/4=2*1/sqrt(2)=sqrt(2)

OpenStudy (anonymous):

is it right? @ikram002p

OpenStudy (anonymous):

or sin pi/4+cos pi/4=1/sqrt(2)+1/sqrt(2)=sqrt(2)

OpenStudy (ikram002p):

@ganeshie8 this is what im thinking off :o but idk if mmm you know mmm identity or hmmm tricky or mmm hmmm m m m mm

ganeshie8 (ganeshie8):

\(\sin \theta - \cos \theta = 0\) gives \(\theta = \frac{\pi}{4} \pm n\pi\) right ?

ganeshie8 (ganeshie8):

cuz period of \(\tan x\) is \(\pi\)

OpenStudy (ikram002p):

yeah also dnt forget 

OpenStudy (anonymous):

3ed?

ganeshie8 (ganeshie8):